# SpotLight: Shadow-Guided Object Relighting via Diffusion

### v1

@@ 4.3 Baselines
We compare SpotLight against recent diffusion-based methods that support light control: DiLightNet [71], IC-Light [73], and Neural Gaffer [30], using their public implementations. While these baselines can relight objects, none are capable of generating a shadow around the object. For fairness, the physically-based renderer Cycles is therefore used to cast a realistic shadow. OpenImageDenoise [1] is also used to denoise the results of all methods. We begin by describing how light control is given to baselines, then proceed to describe per-method specifics.
4.3.1 Light control
In order to ensure fairness in evaluation, we aim to provide each method with a light parametrization that is as close as possible to the ground truth dominant light direction available in the evaluation dataset.
Both DiLightNet [71] and Neural Gaffer [30] require an environment map for lighting. We construct such an environment map with radiance along direction defined as the sum of a spherical gaussian and a constant term
where and are the RGB colors of the light and ambient terms resp., is the dominant light source direction, and is the bandwidth. is obtained by computing the average color over the background image. Panoramas from the Laval Indoor HDR dataset (excluding those in our test set) are used to estimate a single average intensity of the dominant light source, , which is set to the ratio of the integral of the brightest pixels in the panoramas, divided by the integral over all pixels. We further divide this integral by 2 to consider a single hemisphere and avoid overly bright dominant light sources. The light color is defined by , where is the normalized ambient color . We found that the bandwidth parameter did not advantage any specific method and therefore fixed it to , which we typically observe for an indoor light.
4.3.2 Per-method details
DiLightNet [71] DiLightNet requires a prompt describing the object, which we obtain by feeding the ground truth object composited on a black background to BLIP-2 [36] with the prompt “Putting aside the black background, this object is a”. To generate the results, depth and radiance hints are required. We supply DiLightNet with depth information from the 3D model of the object and render the radiance hints using the input environment map (sec. 4.3.1) as outlined in [71]. The rendered object is then composited with the background, along with the rendered shadow.
Neural Gaffer [30] We provide the method with the environment map described in sec. 4.3.1. The ground truth object is composited on a white background before feeding it to the network, as expected by the method. Since Neural Gaffer is trained at lower resolution (), its result is upsampled to then composited on the full resolution background. An erosion using a kernel is applied to the ground truth mask to mitigate boundary artifacts.
IC-Light [73] We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning. We provide a rendering of the object lit by the ground truth environment map, giving a significant advantage to the method. No prompts are used.
@@ 4.4 Experimental results
Tab. 1 presents quantitative results comparing the various baselines (c.f. sec. 4.3) with two versions of SpotLight (c.f. sec. 4.2) on our “reference-based” dataset, where simulated ground truth is available (sec. 4.1). Here, all methods are conditioned on the dominant light direction using their specific light parametrization (sec. 4.3.1). It is worth noting that all three baselines (DiLightNet, IC-Light and Neural Gaffer) are given as input the ground truth object foreground and shadowed background, giving them a significant advantage. Fig. 3 shows qualitative results for each method. We observe that SpotLight with the ZeroComp backbone outperforms the baselines on quantitative metrics and produces visually superior results.
@@
| DiLightNet [71]∗ | 25.36 | 0.950 | 0.059 | 0.024 | 0.044 |
| IC-Light [73]∗ | 28.59 | 0.965 | 0.043 | 0.018 | 0.034 |
| Neural Gaffer [30]∗ | 29.64 | 0.963 | 0.037 | 0.016 | 0.038 |
| SpotLight (RGBX) | 26.44 | 0.956 | 0.052 | 0.018 | 0.041 |
| SpotLight (ZeroComp) | 30.68 | 0.973 | 0.033 | 0.012 | 0.031 |
4.5 User study
A user study was conducted to evaluate the realism of local lighting control. Here, we use the “user-controlled” version of the dataset (see sec. 4.1) to evaluate whether the realism of results obtained by conditioning on a specific lighting direction, even if it is not the dominant light direction.
We evaluate the three methods with the best PSNR from tab. 1, namely SpotLight (with the ZeroComp backbone), Neural Gaffer [30] and IC-Light [73]. Since no ground truth is available, we design a two-alternative forced choice user study, where results obtained by two methods are shown side by side. The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 40 random images are chosen, resulting in 120 pairs of images observed by each participant. We randomly shuffle the 120 pairs for each user, as well as the left-right order. Examples are shown in fig. 5.
The user study was completed by observers. Following Giroux et al. [26], we used the Thurstone Case V Law of Comparative Judgement [60] to obtain a scaling -score for each method, where a higher scale indicates a higher human preference. The results, shown in fig. 4 show that our method is the favorite. The 95% confidence interval error bars also show that the margin in performance is statistically significant, indicating that SpotLight achieves the highest realism.
| Light direction | IC-Light [73] | DiLightNet [71] | Neural Gaffer [30] | Ours (RGBX) | Ours (ZeroComp) |
4.6 Effect of parameter selection
Our method employs parameters that control the degree of shadow guidance ( in eq. 2) and the shadow blending ( in eq. 1). Tab. 2 shows the impact of parameter selection on quantitative metrics, obtained on the “reference-based” dataset where the simulated ground truth is available (c.f. sec. 4.1). Although using no guidance () results in better quantitative performance, we observe qualitatively in fig. 6 that this significantly reduces the visibility of generated shading on the inserted object, thereby reducing the impact of the desired local lighting control. Employing no shadow blending () results in similar metric values but makes shadows much less visible (fig. 7).
| Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
| , (Ours) | 30.68 | 0.974 | 0.033 | 0.012 | 0.031 |
| (no guidance) | 31.65 | 0.976 | 0.030 | 0.011 | 0.029 |
| 28.60 | 0.966 | 0.044 | 0.015 | 0.036 | |
| 30.78 | 0.974 | 0.033 | 0.012 | 0.030 | |
| 29.32 | 0.970 | 0.039 | 0.014 | 0.034 |
@@ 5 Extensions
Shadow control on 2D object composition We demonstrate that SpotLight can be applied to 2D images by leveraging a ZeroComp [75] backbone trained without depth maps, using the normals [68] and albedo [33] estimators. Without access to a 3D model, we cannot rely on a rendering engine to cast shadows. Instead, we employ PixHt-Lab [57] to generate realistic soft shadows from a controllable point light coordinate. Here, a sample without shadow is provided for negative guidance , see sec. 3.3. To estimate the required pixel height of the object cutout, we attribute each pixel -coordinate as the pixel height for demonstration purposes, more sophisticated methods such as [56] could be used. After rendering with the backbone, we enhance the target by adding the high-frequency details from the source image. This is achieved by subtracting a Gaussian-blurred version of the source image ( = 1) from itself (as in unsharp masking) and adding the resulting details to the target. Example 2D compositing results are shown in fig. 8, which demonstrate realistic 2D object relighting from a user-controlled shadow.
| Object | Background | Light-controllable composites | ||
Scribbles as shadows Shadows can be generated in multiple ways: we have demonstrated the use of shadow mapping in 3D and pixel height in 2D. However, one can simply draw the desired shadow! Fig. 9 and SpotLight: Shadow-Guided Object Relighting via Diffusion (first row) show that SpotLight is able to realistically relight the object and refine its shadow even when it is a user-drawn scribble.
@@ 6 Discussion
While the model can realistically estimate relighting from a shadow, generating a shadow map from scratch remains challenging. A promising future direction is to enable end-to-end shadow control using parametric light models, such as point or sphere lights. Additionally, similar approaches could be explored to enable global lighting adjustments for entire images.

### v2

@@ 5.3 Baselines
We compare SpotLight with two types of baselines, that is methods that are conditioned on: the guiding shadow (like SpotLight); or an explicit lighting representation, through a environment map. We now describe per-method details. OpenImageDenoise [1] is also used to denoise the results of all methods to limit the impact of noise generated by diffusion models.
5.3.1 Shadow-conditioned methods
These methods, like ours, use the user-specified guiding shadow as input (obtained with EEVEE, see sec. 5.1).
ZeroComp [78]+SDEdit [44] We create a strong baseline by first running ZeroComp with the shading of the object masked out, and composite the guiding shadow over the background shading intrinsic map. To blend the shadows, we perform a refinement step, inspired by SDEdit [44], by noising this prediction for 50% of the timesteps, and re-running ZeroComp for the remaining timesteps, this time with the shading of the whole image masked out.
IC-Light [76] We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning. We provide a rendering of the object lit by constant ambient lighting as input. No prompts are used.
5.3.2 Light-conditioned methods
Object relighting methods accept an explicit lighting representation as input in the form of a environment map, which we generate using the parametric light direction (see supp.). Unfortunately, these methods focus solely on object relighting—they do not create any cast shadows outside the object. To prevent this absence of shadows from severely hampering the realism of their results, we composite the guiding shadow onto the background image.
DiLightNet [74] DiLightNet requires a prompt describing the object, which we obtain by feeding the object lit by constant ambient lighting and composited on a black background to BLIP-2 [35] with the prompt “Putting aside the black background, this object is a”. DiLightNet also requires depth and radiance hints: we obtain depth from the 3D model of the object and render the radiance hints using the input environment map as outlined in [74]. The rendered object is then composited with the background.
Neural Gaffer [29] The object lit by constant ambient lighting is composited on a white background before feeding it to the network, as expected by the method. Since Neural Gaffer is trained at a lower resolution (), its result is upsampled to then composited on the full resolution background. We noticed Neural Gaffer tends to generate noticeable white boundary artifacts—we remove them through an erosion operation to avoid overly penalizing it.
Guiding shadow | IC-Light [76] | DiLightNet [74] | Neural Gaffer [29] | ZeroComp+SDEdit | Ours | Ground truth |
@@ 5.4 Experimental results
Tab. 1 presents quantitative results comparing the various baselines (c.f. sec. 5.3) with SpotLight (c.f. sec. 5.2) on our “reference-based” dataset, where the ground truth is available (sec. 5.1). Here, all methods are conditioned on the ground truth dominant light direction using their specific light parametrization (sec. 5.3). Fig. 3 shows qualitative results for each method. We observe that SpotLight with the ZeroComp backbone outperforms the baselines on quantitative metrics and produces visually superior results.
(a) Overall realism study
(b) Lighting control study
5.5 User studies
Two user studies were conducted to assess the perceptual realism and controllability of SpotLight compared to baselines. In both cases, the Thurstone Case V Law of Comparative Judgement [61] is used to obtain a scaling -score for each method, where a higher scale indicates a higher human preference. Here, we give a high-level overview of those user studies. Refer to the supplementary material for additionnal details and studies.
User study I: Overall realism
Here, we use the “user-controlled” version of the dataset (see sec. 5.1) to evaluate the realism of results obtained by conditioning on a specific lighting direction, even if it is not aligned with the scene lighting conditions. We compare SpotLight against the three baselines with the best PSNR from tab. 1, namely Neural Gaffer [29], IC-Light [76], and ZeroComp [78]+SDEdit [44]. Since no ground truth is available, we design a two-alternative forced choice user study, where results obtained by two methods are shown side by side. The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 20 random images are chosen (we ensure images where shadows are clearly visible are selected), resulting in 120 pairs of images observed by each participant. We randomly shuffle the 120 pairs for each user, as well as the left-right order. Note that we also add three sentinel images to discard observers that did not understand the task. Examples are shown in fig. 3.
User study I was completed by observers. The results, shown in fig. 4a, show that our method is the favorite, indicating that SpotLight achieves the highest realism, with a statistically significant difference to Neural Gaffer and IC-Light.
User study II: lighting control Since the goal of our method is to give full artistic control over the object relighting, we design a second two-alternative forced choice user study to evaluate lighting controllability separately from the realism. The users were shown two videos side by side and were asked to “Select the video where the light direction has the most impact on the lighting of the object, by considering both the appearance of the object and its shadows.”. The videos show a virtual object relit by a rotating light source (see supp.). Here, we compared the two methods with the highest realism according to user study I: SpotLight and ZeroComp+SDEdit. Since this task aims to evaluate the local relighting within the object mask only, we replace the shadowed background with the output from our method.
User study II was completed by observers. As seen in fig. 4b, SpotLight achieves a statistically significant improvement in performance over ZeroComp+SDEdit, despite the low number of participants, which we attribute to the use of our local guidance strategy.
5.6 Effect of parameter selection
Our method employs parameters that control the degree of shadow guidance ( in eq. 3) and the shadow blending ( in eq. 2). Tab. 2 shows the impact of parameter selection on quantitative metrics, obtained on the “reference-based” dataset where the ground truth is available (c.f. sec. 5.1). Although using no guidance () results in better quantitative performance, we observe qualitatively in fig. 5 that this significantly reduces the visibility of generated shading on the inserted object, thereby reducing the impact of the desired local lighting control. Employing no shadow blending () results in similar metric values but makes shadows much less visible (refer to the supp. for a qualitative comparison).
@@ 5.6 Effect of parameter selection
, (Ours) | 30.68 | 0.974 | 0.033 | 0.012 | 0.030 |
(no guidance) | 31.69 | 0.976 | 0.029 | 0.011 | 0.029 |
28.68 | 0.966 | 0.043 | 0.015 | 0.036 | |
30.81 | 0.974 | 0.032 | 0.012 | 0.029 | |
29.24 | 0.969 | 0.039 | 0.014 | 0.034 |
@@ 6 Applications
Scribbles as shadows Shadows can be generated in multiple ways: we have demonstrated the use of shadow mapping in 3D and, in the supp., pixel height in 2D [58]. However, one can simply draw the desired shadow! Fig. 6 shows that SpotLight is able to realistically relight the object and refine its shadow even when it is a user-drawn scribble.
Full image relighting As shown in fig. 7, we can extend SpotLight to achieve full-scene relighting. Although our method is not explicitly designed for this task, we find our shadow guidance to be effective in plausibly relighting the background. First, we relight the object using SpotLight with the desired guiding shadow. Next, we apply a version of ZeroComp trained with an inverted shading mask, using only the object shading to relight the background. We repeat this process for different light directions and average the resulting predictions to obtain an average relit background. Finally, we compute a per-pixel highlight map by dividing each background prediction by the average and multiplying it by the original background. This process ensures that background shadows and highlights consistently align with the object’s shadow.
@@ 7 Discussion
Incorrect user-defined shadow | Correct shadow |
Limitation: physically incorrect shadows Since SpotLight imposes no constraints on the input shadow, the user is free to specify a shadow direction that potentially contradicts the real lighting direction in the background image. In this case, SpotLight will still attempt to generate an image that looks as realistic as possible, given the shadow conditioning, as shown in fig. 8.
@@
While the model can realistically generate relighting results from a shadow and despite showing that this can be done in multiple ways (shadow mapping, 2D generation, hand-drawn scribbles), the process of generating a shadow from scratch could prove challenging to some. A promising future direction is to enable end-to-end shadow control using parametric light models, such as point or sphere lights. Additionally, similar approaches could be explored to enable global lighting adjustments for entire images.